OpenStudy (anonymous):

Write as a single fraction: (6x+4y/9x) - (6x-7y/3x)+2

OpenStudy (muzzack):

Step by step solution : Step 1 : 7xy Simplify 6x - ——— 3 Rewriting the whole as an Equivalent Fraction : 1.1 Subtracting a fraction from a whole Rewrite the whole as a fraction using 3 as the denominator : 6x 6x • 3 6x = —— = —————— 1 3 Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Adding fractions that have a common denominator : 1.2 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: 6x • 3 - (7xy) 18x - 7xy —————————————— = ————————— 3 3 Equation at the end of step 1 : y (18x-7xy) ((6x+((4•—)•x))-—————————)+2 9 3 Step 2 : 4xy Simplify 6x + ——— 9 Rewriting the whole as an Equivalent Fraction : 2.1 Adding a fraction to a whole Rewrite the whole as a fraction using 9 as the denominator : 6x 6x • 9 6x = —— = —————— 1 9 Adding fractions that have a common denominator : 2.2 Adding up the two equivalent fractions 6x • 9 + 4xy 4xy + 54x ———————————— = ————————— 9 9 Equation at the end of step 2 : (4xy + 54x) (18x - 7xy) (——————————— - ———————————) + 2 9 3 Step 3 : 4xy+54x 18x-7xy Simplify ——————— - ——————— 9 3 Pulling out like terms : 3.1 Pull out like factors : 4xy + 54x = 2x • (2y + 27) Pulling out like terms : 3.2 Pull out like factors : 18x - 7xy = -x • (7y - 18) Calculating the Least Common Multiplier : 3.3 Find the Least Common Multiple (L.C.M) The left denominator is : 9 The right denominator is : 3 Factor the left and right denominators, counting the number of times each factor appears : |———————————————|—————————————————————————————————————| | | Number of times each prime factor | | | appears the factorization of: | | | | | Prime | Left Right L.C.M = Max | | Factors |Denominator Denominator {Left,Right}| | —————————— |——————————— ——————————— ————————————| | 3 | 2 1 2 | |———————————————|—————————————————————————————————————| |Product of all | | | Prime Factors | 9 3 9 | |———————————————|—————————————————————————————————————| Least Common Multiple: 9 Calculating Multipliers : Calculate multipliers for the two fractions L.C.M L. Multiplier = —————————————— = 1 L. Denominator L.C.M R. Multiplier = —————————————— = 3 R. Denominator Making Equivalent Fractions : 3.5 Rewrite the two fractions into equivalent fractions Two fractions are called equivalent if they have the same numeric value. For example : 1/2 and 2/4 are equivalent, y/(y+1)2 and (y2+y)/(y+1)3 are equivalent as well. To calculate equivalent fraction , multiply the Numerator of each fraction, by its respective Multiplier. L. Mult. • L. Num. 2x • (2y+27) —————————————————— = ———————————— L.C.M 9 R. Mult. • R. Num. -x • (7y-18) • 3 —————————————————— = ———————————————— L.C.M 9 Adding fractions that have a common denominator : 3.6 Adding up the two equivalent fractions 2x • (2y+27) - (-x • (7y-18) • 3) 25xy ————————————————————————————————— = ———— 9 9 Equation at the end of step 3 : 25xy ———— + 2 9 Step 4 : 25xy Simplify ———— + 2 9 Rewriting the whole as an Equivalent Fraction : 4.1 Adding a whole to a fraction Rewrite the whole as a fraction using 9 as the denominator : 2 2 • 9 2 = — = ————— 1 9 Adding fractions that have a common denominator : 4.2 Adding up the two equivalent fractions 25xy + 2 • 9 25xy + 18 ———————————— = ————————— 9 9 Final result : 25xy + 18 ————————— 9

geerky42 (geerky42):

Did you copy it from somewhere on internet? lol... @Muzzack

OpenStudy (muzzack):

nope

geerky42 (geerky42):

haha ok I "believe" you. Does this help you though? @Calvinvd